The Perimeter of a rectangle is 24 inches.Find the dimensions if its length is 3 inches greater than its widths.
Found 2 solutions by Annabelle1, acchs_cheerleader:
Answer by Annabelle1(69)

let its width be x
so the length will then be x+3 (3inches greater)
The perimeter is then equal to
x+x+(x+3)+(x+3)
this is equal to 24
x+x+(x+3)+(x+3)=24
4x+6=24
4x+6-6=24-6 (subtracting 6 from both sides of the equation)
4x=18
x=18/4
x=4.5
so the width is 4.5
the length will be 4.5+3=7.5



Answer by acchs_cheerleader(6)

The formula for the perimeter of a rectangle is 2(l+w)=perimeter. To find the answer you do:
1) 2(l+w)=P
2) 2(3+w+w)=24 [ since the length is 3 inch greater it means add ]
3) 2(3+2w)=24 [ you find the like terms and add ]
4) 6+4w=24 [ you distribute the 2 by multiplying]
5) 4w=18 [ you subtract 6 to both sides ]
6) w = 4.5 [ you divide 4 to both sides ]
________________________________________________________________________________
CHECK:
1) 2(l+w)=P
2) 2(3+4.5+4.5)=24
3) 2(12)=24
4) 24=24